If X greater than 0 and Y greater than 0 and X+Y=18 then the minimum value of XY is
A.   -1
B.    1
C.    2
D.   -2

Grade:12

1 Answers

Arun
3 years ago
Dear student
 
Let A = xy
A = x (18-x)
 
dA/dx = 18 -2x = 0
x = 9
 
Now d²A /dx² = -2 (maximum)
 
Hence maximum of A will exist when x = y = 9
 
Hence A = xy = 9*9 = 81